OpenStudy (anonymous):

solve the rational equation 3x/4 = 5x + 1/3???

OpenStudy (anonymous):

\[\frac{3x}{4}=\frac{5x+1}{3}\]?

OpenStudy (anonymous):

Do I begin by multiplying 3 to 3x and 4 to 5x + 1? would it become 9x/12x = 20x + 4/12

OpenStudy (anonymous):

yes, you "cross multiply" and get \[3\times 3x=4(5x+1)\]

OpenStudy (anonymous):

so then its 9x = 20x + 1? and wat do i do after?

OpenStudy (anonymous):

don't forget the distributive law !!

OpenStudy (anonymous):

wait, what would it be then?

OpenStudy (anonymous):

\[9x=20x+4\]

OpenStudy (anonymous):

oh yeah! ok. so then i subtract 4 or divide by 9?

OpenStudy (anonymous):

probably subtract \(20x\) from both sides then divide by \(-11\)

OpenStudy (anonymous):

r u sure?

OpenStudy (anonymous):

pretty sure, yes

OpenStudy (anonymous):

\[9x=20x+4\] \[-11x=4\] \[x=-\frac{4}{11}\]

OpenStudy (anonymous):

alright. thanx for the help :)

OpenStudy (anonymous):

yw